A 0.006765 mol sample of an organic compound was burned in oxygen in a bomb calorimeter. The temperature of the calorimeter increased from 25.1 °C to 29.9 °C. If the heat capacity of the calorimeter is 4.78 kJ (°C)−1, then what is the constant volume heat of combustion of this compound, in kilojoules per mole? (Remember to include a "+" or "-" sign in your answer, as appropriate.)

Enter your answer with 3 significant figures. Do not include units as part of your answer.
1
Expert's answer
2018-03-07T11:58:07-0500
Find the heat emitted after burning:
Q = C × (T2-T1) = 4.78 × (29.9-25.1) = +22.944 (kJ)
Find the constant volume heat of combustion per mole:
q = Q/ν = (+22.944)/(0.006765) = +3391.6 (kJ/mol)